OpenStudy (anonymous):

Find the first six terms of the sequence. a1 = -8, an = 5 • an-1

OpenStudy (anonymous):

Explanation please!

OpenStudy (mathmale):

Hello, Tonio, Looks like the first term of your sequence is -8 (which you have called "a1"). The nth term is 5*a(n-1) (which you have called an-1). Let me type that in Equation editor:\[a _{1}=-8;a _{n}=5*a _{n-1}\]

OpenStudy (mathmale):

So, if the first term, a1=-8, the second term is \[a _{2}=5a _{2-1}=5a _{1}=5(-8)=-40\]

OpenStudy (mathmale):

Please look that over. Then see whether you can find the third, fourth, fifth and 6th terms. Please write out your work (here, or on paper and then share the image with me).

OpenStudy (anonymous):

oh okay so \[a _{3}=5a _{3-2}=5_{a2}=5(-40)=-200\]?

OpenStudy (anonymous):

@mathmale

OpenStudy (mathmale):

Looks good. Continue until you have a total of 6 terms.

OpenStudy (anonymous):

so the terms are -8;-40;-1,000;-50,000;-250,000

OpenStudy (anonymous):

5000 , 25000 sorry

OpenStudy (mathmale):

Tonio: what's 5(-40)?

OpenStudy (mathmale):

You're on the right track, but have a few corrections to make.

OpenStudy (anonymous):

oh I forgot -200

OpenStudy (anonymous):

I have it on my paper lol oops

OpenStudy (mathmale):

Are you satisfied that you understand what to do, or would you like more discussion?

OpenStudy (anonymous):

I understand some of it, let me find another equation really quick.

OpenStudy (mathmale):

OK

OpenStudy (anonymous):

Find an equation for the nth term of the arithmetic sequence. -3, -5, -7, -9, ...

OpenStudy (anonymous):

@mathmale

OpenStudy (mathmale):

How do you get -5 from -3? Then: how to you get -7 from -5? Try to figure out what happens to produce a new term from the preceding term.

OpenStudy (anonymous):

well you minus 2

OpenStudy (mathmale):

Yes, you subtract 2 from one term to obtain the next.

OpenStudy (mathmale):

\[a _{n+1}=a _{n}-2\] Try testing this formula. Does it correctly predict each new a(n) value?

OpenStudy (anonymous):

so \[a _{n}=-3+-2(n) \]

OpenStudy (mathmale):

Best answer I could give you at this point is to ask you to do the same thing: Test whether or not your result (formula) correctly predicts each new term.

OpenStudy (anonymous):

okay so your saying to plug in numbers?

OpenStudy (anonymous):

when you say \[a _{n+1}\] does this mean to add the last answer you just had or insert the answer just got?

OpenStudy (mathmale):

Look at the sequence you've given me: -3, -5, -7, -9, ... What is a1? what is a2? How do we get a2 from a1? Test your own formula: try predicting a2 when a1--3. Either your formula works or it does not. Simiilarly, try testing my formula; again, either it works or it does not In this manner you can determine the correct predictive / recursive formula.

OpenStudy (mathmale):

when you say an+1 does this mean to add the last answer you just had or insert the answer just got? \[a _{n+1}\] represents the next term in the sequence when you start with term \[a _{n}\]

OpenStudy (anonymous):

well I already know my formula does not work.

OpenStudy (mathmale):

In this case your first term, a1, is -3, and your n=1. You want to find the 2nd term, a2, from a1.

OpenStudy (mathmale):

But in deciding that, you learned something of value, right? It's very important to know how to check a recursive formula, as well as to actually do it!

OpenStudy (anonymous):

yeah learing something I hate online school teachers do not help at all, but the only equation that would fit in this an=-3-2

OpenStudy (anonymous):

learning*

OpenStudy (mathmale):

So, where are we now? Do we have a recursion formula that "works"?

OpenStudy (anonymous):

well it will work, but does the equation supposed have a n in it?

OpenStudy (mathmale):

Yes, it does, but the n is a subscript, not a multiplier.

OpenStudy (mathmale):

\[a _{n+1}=a _{n}-2\] Your first term is a1=-3. Please predict a2.

OpenStudy (mathmale):

a1-2=-3-2=??

OpenStudy (anonymous):

Okay so it would equal -5

OpenStudy (anonymous):

here let me show you my answer

OpenStudy (mathmale):

Is that what you'd hoped for, or not?

OpenStudy (anonymous):

an = -3 - 2 an = -3 + -2(n) an = -3 + -2(n + 1) an = -3 + -2(n - 1)

OpenStudy (anonymous):

yeah but none of these look like your equation

OpenStudy (mathmale):

(Thinking)

OpenStudy (anonymous):

I was thinking C but not really

OpenStudy (mathmale):

Again, I believe you'd learn the most by actually substituting n=1,2,3,4, etc., in the recursion equation (C). Does the equation correctly predict a(n) in each case n=1,2,3, 4, etc.??

OpenStudy (mathmale):

It appears that there may be more than 1 correct answer to this problem; one from the list given to you , one from the recursion formulas that you and I created.

OpenStudy (anonymous):

yeah that is why I am so confused

OpenStudy (mathmale):

Too bad you apparently MUST choose one of the four options given; that squelches creativity in finding other possible approaches.

OpenStudy (mathmale):

Don't be offended, but I'm going to ask you to go back and double check to ensure that you ve copied the four answer options correctly.

OpenStudy (anonymous):

ok

OpenStudy (mathmale):

Checking?? While waiting, I've reviewed the problem and have come up with a recursion formula that seems to work as it's supposed to:

OpenStudy (mathmale):

\[a _{n}=-3-2(n-1)\]

OpenStudy (mathmale):

I'd suggest you check it. I see you're viewing someone else's question. do let me know when you're back if you'd still like to finish up the work on your own question.